OpenStudy (anonymous):

The average speed of one airplane is 200 mph faster than that of an automobile. If the automobile starts at 3 a.m. and the airplane starts from the same place at 11 a.m., they will have traveled the same number of miles by 1:00 p.m. If r represents the rate of the automobile, how many miles does the plane travel? 200r 2(200 + r) 10r

OpenStudy (perl):

lets make an equation for the car and the airplane. distance = rate * time

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

How?

OpenStudy (perl):

How about this. Let d1 = distance car travels r1 = rate of car, t1 = number of hours the car travels. d2 = distance plane travels r2 = rate of plane t2 = number of hours plane travels Therefore we have car equation: d1 = r1 * t1 plane equation d2 = r2 * t2

OpenStudy (perl):

d1 stands for a subscript $$ d_1 $$

OpenStudy (perl):

we are given two pieces of information, r2 = r1 + 200 and that d1 = d2 at 1 P.M.

OpenStudy (perl):

At 1 PM the car has traveled 10 hours, since it started driving at 3 A.M. At 1 PM the plane has traveled 2 hours, since it started flying at 11 AM. Therefore we have : d1 = r1 * t1 d2 = (200 + r1) * t2 and substituting time t1 = 10, and t2 = 2, d1 = d2 r1 * (10) = ( 200 + r1) * 2

OpenStudy (anonymous):

So it's 2(200 + r) ..?

OpenStudy (perl):

correct , since r1 is the same as r in the directions

OpenStudy (perl):

but d1 = 10r , so it is also 10 r

OpenStudy (anonymous):

Thanks so much for explaining it to me:)

OpenStudy (perl):

also you can solve that equation for r , you might as well. 10 r = 2 ( 200 + r) 10r = 2*200 + 2r 8r = 400 r = 400 / 8 = 50 mph so the plane went 250 mph , since it is 200 mph faster than the car rate r. and you can see that 10 * 50 = 500 miles 2 ( 200 + 50) = 2 * 250 = 500 miles so it works :)

OpenStudy (anonymous):

Yes it sure does..! :) I appreciate all the help. I might need some more help in a little while. Would you mind assisting me again?

OpenStudy (perl):

sure

OpenStudy (anonymous):

Okie dokie! I'll let you know